# Expert-Guided Diffusion Planner for Auto-bidding

### v1

@@ Abstract.
Auto-bidding is extensively applied in advertising systems, serving a multitude of advertisers. Generative bidding is gradually gaining traction due to its robust planning capabilities and generalizability. In contrast to traditional reinforcement learning-based bidding, generative bidding does not rely on the Markov Decision Process (MDP) exhibiting superior planning capabilities in long-horizon scenarios. Conditional diffusion modeling approaches have demonstrated significant potential in the realm of auto-bidding. However, relying solely on return as the optimality condition is weak to guarantee the generation of genuinely optimal decision sequences, lacking personalized structural information. Moreover, diffusion models’ t-step autoregressive generation mechanism inherently carries timeliness risks. To address these issues, we propose a novel conditional diffusion modeling method based on expert trajectory guidance combined with a skip-step sampling strategy to enhance generation efficiency. We have validated the effectiveness of this approach through extensive offline experiments and achieved statistically significant results in online A/B testing, achieving an increase of 11.29% in conversion and a 12.35% in revenue compared with the baseline.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
The bidding mechanism of online advertising platforms, as a core component of modern digital marketing (Yuan et al., 2013; Wang et al., 2017; Adikari and Dutta, 2015; Evans, 2009), remains pivotal in enabling businesses to target audiences and drive commercial outcomes precisely. However, the inherent volatility of dynamic market conditions and the intricate diversity of user behaviors, combined with the need to process trillions of daily ad impressions in real time (Guo et al., 2024a; Johnson et al., 2020), expose dual challenges for traditional manual bidding strategies. Advertisers must constantly modify their bid prices to keep up with the constantly changing competitive bidding environment, which requires extensive knowledge in the field. Still, they frequently do not reach the best results.
To address these challenges, Ad platforms have developed auto-bidding services (Wen et al., 2022; Aggarwal et al., 2024; Lin et al., 2024) to simplify bid optimization for advertisers. Through carefully designed bidding strategies that account for various advertising factors (such as impression distributions, budgets, and cost constraints), these services automatically set bid prices for each impression opportunity. Various advanced technologies have been recently adopted to meet these demands, including Reinforcement Learning (RL)-based and Generative methods.
@@
Reinforcement learning models the bidding process as a Markov Decision Process (MDP) and has been widely adopted in automated bidding systems, demonstrating consistent performance enhancements across diverse advertising scenarios (He et al., 2021; Lin et al., 2023; Mou et al., 2022). However, conventional reinforcement learning approaches grounded in Markov conditional independence assumptions and dynamic programming principles exhibit inherent limitations in bidding effectiveness, which primarily stems from (1) severe cumulative error propagation due to extended decision horizons in bidding sequences and (2) ignoring the impact of long-term dependencies in the bidding environment (Zhu et al., 2023).
Generative bidding demonstrates superior long-horizon planning capabilities and has gained significant traction.(Li et al., 2024) applies Decision Transformer to conduct sequential planning. (Guo et al., 2024b) applies a diffusion planner, incorporating influences from preceding states and enabling extended-horizon strategic planning.
However, applying conditional diffusion models to bidding faces two key challenges. First, existing approaches condition cumulative returns as a single scalar hyperparameter during inference, leading to weak guidance due to (i) lack of personalization and (ii) excessively large solution space satisfying the scalar constraint, thereby hindering convergence to optimal solutions. Second, the online bidding module’s strict latency requirements conflict with the T-step iterative sampling process inherent to diffusion models.
@@
To tackle the challenges previously discussed, we introduce a Expert-Guided Diffusion planner for auto-Bidding (EGDB). Firstly, we propose a blended-forcing mechanism that synergistically integrates teacher-forcing stabilization and VAE-enhanced decode-forcing. It enables precise imitation of expert trajectories while proactively exploring latent optimal bidding patterns beyond observed demonstrations. Secondly, building upon the refined expert guidance, we devise a dual-conditioned planner that implicitly incorporates expert behavioral semantics while explicitly enforcing domain constraints (e.g., budget), achieving principled adaptation to dynamic auction environments. Finally, to overcome the inherent latency of iterative diffusion, we develop a step-skipping strategy that reduces inference steps without compromising action generation quality - a crucial advancement for real-time bidding systems. We demonstrate how this framework achieves state-of-the-art offline and online performance through comprehensive experiments.
In conclusion, our contributions are depicted as follows:
@@
  • We explicitly construct expert demonstrations and leverage them to provide personalized guidance for the diffusion generation process.
@@
  • To ensure training-inference consistency, we introduce a Variational Auto-Encoder (VAE) to model expert inherent behaviors, coupled with blended-forcing techniques to mitigate error accumulation and enhance training efficiency.
@@ 2. Related Work
Offline Reinforcement Learning. In recent years, Offline reinforcement learning (Offline RL) has emerged as a pivotal paradigm for decision-making. It aims to learn optimal policies solely from static datasets of historical interactions, without requiring online environment exploration. This paradigm addresses scenarios where allowing partially trained agents to interact with the environment could lead to safety risks or substantial costs. Offline RL methods (Kostrikov et al., 2022; Fujimoto et al., 2019; Kumar et al., 2020; Chen et al., 2021) address two critical challenges—distribution shift and extrapolation errors-through distinct methodological innovations. Implicit Q-Learning (IQL) (Kostrikov et al., 2022) tackles distribution shift by implicitly estimating optimal actions through upper expectile regression on state values, suppressing extrapolation errors via asymmetric Q-updates that prioritize in-distribution high-value actions. Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) combines a variational autoencoder (VAE) with Q-filtering to restrict policy updates to data-supported regions, minimizing overestimation risks. Conservative Q-Learning (CQL) (Kumar et al., 2020) explicitly regularizes Q-values with theoretical guarantees, penalizing out-of-distribution (OOD) actions through conservative value bounds. In contrast, Decision Transformer (DT) (Chen et al., 2021) circumvents value functions entirely by reframing RL as sequence modeling, autoregressively generating actions conditioned on historical trajectories to eliminate bootstrapping errors and OOD queries.
@@
Diffusion Models. Through their unique stochastic denoising mechanisms, diffusion models (Ho et al., 2020; Song et al., 2020) have established themselves as a dominant paradigm in image synthesis, capable of generating high-fidelity data via Markov chain-based iterative refinement. These models can be divided into two categories: conditional diffusion and unconditional diffusion. Conditional diffusion models (Fu et al., 2024) incorporate additional information (e.g., class labels, text prompts, or environmental states) to steer the generation process, enabling targeted synthesis. While unconditional diffusion models (Heng et al., 2024) generate data from pure noise without external guidance, excelling in capturing the intrinsic distribution of the training data. Owing to the powerful capabilities of diffusion models, they have been widely applied across various fields, particularly in decision-making. (Hansen-Estruch et al., 2023) proposes IDQL, a reinforcement learning framework that integrates implicit Q-learning with diffusion policies in an actor-critic architecture. (Wang et al., 2022) leverages diffusion models to capture complex, multimodal behavior distributions from static datasets. (Chen et al., 2022) uses diffusion models to more accurately fit the complex distribution of behavioral policies and avoids selecting unseen actions that are out of the data distribution during training.
@@
Auto-bidding. Auto-bidding, which is primarily used in online advertising auction scenarios, dynamically adjusts advertisers’ bidding strategies to maximize advertising effectiveness (such as clicks, conversions, or sales) while meeting constraints like budgets and return on ad spend (ROAS). RL-based methods are widely used in auto-bidding because they dynamically optimize bids in complex advertising environments and balance exploration and exploitation. (Cai et al., 2017) proposes a reinforcement learning-based framework to optimize real-time bidding strategies in display advertising. (He et al., 2021) proposes a generalized framework for optimizing real-time bidding strategies under diverse budget and key performance indicator (KPI) constraints, dynamically adjusting bid parameters via reinforcement learning to maximize campaign value while ensuring compliance with advertiser requirements. (Jin et al., 2018) explores multi-agent reinforcement learning for real-time bidding in display advertising. Recent studies (Gao et al., 2025; Guo et al., 2024b) have increasingly explored generative auto-bidding strategies in online advertising. (Gao et al., 2025) leverages Decision Transformer (Chen et al., 2021) to model sequential dependencies and historical auction dynamics, enabling adaptive bidding strategy optimization by autoregressively generating context-aware bid adjustments aligned with long-term advertiser objectives. (Guo et al., 2024b) directly models the correlation between the return and the entire trajectory and uses a conditional diffusion for bid generation.
@@ 3. Preliminary
In this section, we outlined the general bidding and auction tasks framework and examined the architecture of generative diffusion models. To prevent any confusion regarding subscripts in the subsequent discussion, it should be noted that $k$ denotes the diffusion timesteps, while $t$ represents the position of the trajectory $\tau$.
3.1. Formulation of Bidding Problem
@@ 3.2. Diffusion Model
3.3. Classifier-free Diffusion Guidance
@@
(2) $\max_{\theta}\mathbb{E}_{\tau\sim D}\left[\log p_{\theta}\left(x_{0}(\tau)\mid y(\tau)\right)\right]$
@@
where $q$ represents the forward process in which noises are gradually added to the trajectory, while $p_{\theta}$ is the reverse process where a model is used for denoising.
We model the forward process $q\left(x_{k+1}(\tau)\mid x_{k}(\tau)\right)$ via diffusion over states, where:
@@
where $s_{t}$ is a one-dimensional vector, including real-time bidding context, and $x_{k}(\tau)$ is a noise sequence of states which can be represented by a two-dimensional array where the first dimension is the time periods and the second dimension is the state values. Given $x_{k}(\tau)$, the forward process of the diffusion model can be represented as follows:
(5) $q\left(x_{k}(\tau)\mid x_{k-1}(\tau)\right)=\mathcal{N}\left(x_{k}(\tau);\sqrt{1-\beta_{k}}x_{k-1}(\tau),\beta_{k}\mathbf{I}\right),$
@@
When serving at period $t$, we first sample from an initial trajectory $x^{\prime}_{k}(\tau)\sim\mathcal{N}(0,\mathbf{I})$ and assigns the history states $s_{0:t}$ into it. Then, we can sample predicted states with the reverse process recursively by
(8) $x^{\prime}_{k-1}(\tau)=\mu_{\theta}\left(x^{\prime}_{k}(\tau),y(\tau),k\right)+\sqrt{\beta_{k}}z,$
@@ 4. Methodology
In this section, we present a comprehensive description of the proposed Expert-Guided Diffusion planner for auto-Bidding ( EGDB ) framework. We begin by formalizing the core mechanism of Blended Forcing with Optimal Expert in Section 4.1. Next, in Section 4.2, we elaborate on the Expert-guided Conditional Planner that translates blended signals into structured action sequences. Finally, Section 4.3 details the Efficient Action Generation module designed for real-time execution. The overall architecture is illustrated in Figure 1.
4.1. Blended Forcing with Optimal Expert
To address the bidding problem outlined in Equation 1, we initially aim to generate expert trajectories. To concentrate on a single bidding agent, we remove the subscript $i$, assuming all other agents are fixed. We transform the linear programming (LP) function into its dual form. Assuming the auction is truthful, the bidding formula will produce an outcome corresponding to the optimal primal solution $b_{j}$ (Aggarwal et al., 2024). For each query $j$:
(10) $\mathbf{x}^{e}(\tau)=b_{j}:=\frac{1+\alpha_{c}\cdot C_{i}}{\alpha_{b}+\alpha_{c}}\cdot v_{j},$
where $\alpha_{b}$ represents the influence of budget constraints and $\alpha_{c}$ denotes the impact of constraints of cost.
Then, we can refine the $\mathbf{x}^{e}(\tau)$ to fully capture the latent expert behaviors through the blended-forcing mechanism in the training phase. As shown in Figure 1, it comprises two operations: (a) the teacher-forcing technique ($TF(\cdot)$) to imitate the expert behaviors more stably, and (b) the decode-forcing mechanism ($DF(\cdot)$) to explore more optimal expert behaviors through an extra Variational Auto-Encoder(Kingma, 2013) (VAE).
Teacher Forcing (TF), initially introduced in Recurrent Neural Networks (RNNs) (Williams and Zipser, 1989), is a training technique where the ground truth sequence is fed as input to the model at each timestep, rather than using its previous predictions. This approach effectively mitigates error accumulation in autoregressive models by aligning training dynamics with the true data distribution. However, a critical limitation arises during inference: the absence of ground truth forces the model to rely entirely on its own (potentially erroneous) predictions, leading to a mismatch between training and inference input distributions. This phenomenon, termed Exposure Bias, often results in degraded performance in real-world scenarios. Inspired by this technique, we introduce TF and use expert trajectories serving as ground truth sequences to help the model imitate expert behaviors.

### v2

@@ Abstract.
Auto-bidding is widely used in advertising systems, serving a diverse range of advertisers. Generative bidding is increasingly gaining traction due to its strong planning capabilities and generalizability. Unlike traditional reinforcement learning-based bidding, generative bidding does not depend on the Markov Decision Process (MDP), thereby exhibiting superior planning performance in long-horizon scenarios. Conditional diffusion modeling approaches have shown significant promise in the field of auto-bidding. However, relying solely on return as the optimality criterion is insufficient to guarantee the generation of truly optimal decision sequences, as it lacks personalized structural information. Moreover, the auto-regressive generation mechanism of diffusion models inherently introduces timeliness risks. To address these challenges, we introduce a novel conditional diffusion modeling approach that integrates expert trajectory guidance with a skip-step sampling strategy to improve generation efficiency. The efficacy of this method has been demonstrated through comprehensive offline experiments and further substantiated by statistically significant outcomes in online A/B testing, yielding an 11.29% increase in conversions and a 12.36% growth in revenue relative to the baseline.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ‘25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761497isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommender systems
@@ 1. Introduction
The bidding mechanism of online advertising platforms, as a core component of modern digital marketing (Yuan et al., 2013; Wang et al., 2017; Adikari and Dutta, 2015; Evans, 2009), remains pivotal in enabling businesses to target audiences and drive commercial outcomes precisely. However, the inherent volatility of dynamic market conditions and the intricate diversity of user behaviors, combined with the need to process trillions of daily ad impressions in real time (Guo et al., 2024a; Johnson et al., 2020), expose dual challenges to traditional manual bidding strategies. Advertisers must continually adjust their bid prices to stay competitive in a rapidly evolving bidding environment, which demands extensive knowledge in the field. Still, they frequently do not reach the best results.
@@
To address these challenges, advertising platforms have developed auto-bidding services (Wen et al., 2022; Aggarwal et al., 2024; Lin et al., 2024) to simplify bid optimization for advertisers. Through carefully designed bidding strategies that account for various advertising factors (such as impression distributions, budgets, and cost constraints), these services automatically set bid prices for each impression opportunity. Various advanced technologies have recently been adopted to meet these demands, including Reinforcement Learning (RL)-based and generative methods. Reinforcement learning models the bidding process as a Markov Decision Process (MDP) and has been widely adopted in automated bidding systems, demonstrating consistent performance enhancements across diverse advertising scenarios (He et al., 2021; Lin et al., 2023; Mou et al., 2022). However, conventional reinforcement learning approaches grounded in Markov conditional independence assumptions and dynamic programming principles exhibit inherent limitations in bidding effectiveness, mainly stemming from (i) severe cumulative error propagation due to extended decision horizons in bidding sequences and (ii) ignoring the impact of long-term dependencies in the bidding environment (Zhu et al., 2023). Generative bidding demonstrates superior long-term planning capabilities and has gained significant traction.(Li et al., 2024) applies the Decision Transformer to conduct sequential planning. (Guo et al., 2024b) applies a diffusion planner, incorporating influences from preceding states and enabling extended-horizon strategic planning.
@@
However, applying conditional diffusion models to bidding faces two key challenges. First, existing approaches condition cumulative returns as a single scalar hyperparameter during inference, leading to weak guidance due to (i) lack of personalization and (ii) vast solution space satisfying the scalar constraint, thereby hindering convergence to optimal solutions. Second, the online bidding module’s strict latency requirements conflict with the T-step iterative sampling process inherent to diffusion models. To tackle these challenges discussed previously, we introduce a novel framework called Expert-Guided Diffusion planner for Auto-Bidding (EGDB). Firstly, we propose a blended-forcing mechanism that synergistically integrates teacher-forcing stabilization and VAE-enhanced decode forcing. It enables precise imitation of expert trajectories while proactively exploring latent optimal bidding patterns beyond observed demonstrations. Secondly, building upon the refined expert guidance, we devise a dual-conditioned planner that implicitly incorporates expert behavioral semantics while explicitly enforcing domain constraints (i.e., budgets), achieving principled adaptation to dynamic auction environments. Finally, to overcome the inherent latency of iterative diffusion, we develop a step-skipping strategy that reduces inference steps without compromising the quality of action generation — a crucial advancement for real-time bidding systems. We demonstrate how this framework achieves state-of-the-art offline and online performance through comprehensive experiments.
In conclusion, our main contributions can be summarized in the following four parts:
@@
  • We construct expert demonstrations explicitly and leverage them to provide personalized guidance for the diffusion generation process.
@@
  • To ensure training-inference consistency, we introduce a Variational Auto-Encoder to model expert inherent behaviors, coupled with blended-forcing techniques to mitigate error accumulation and enhance training efficiency.
@@ 2.1. Offline Reinforcement Learning
In recent years, Offline reinforcement learning (RL) has emerged as a pivotal paradigm for decision-making. It aims to learn optimal policies solely from static datasets of historical interactions, without requiring online environment exploration. This paradigm addresses scenarios where allowing partially trained agents to interact with the environment could lead to safety risks or substantial costs. Offline RL methods (Kostrikov et al., 2022; Fujimoto et al., 2019; Kumar et al., 2020; Chen et al., 2021) address two critical challenges—distribution shift and extrapolation errors-through distinct methodological innovations. Implicit Q-Learning (IQL) (Kostrikov et al., 2022) tackles distribution shift by implicitly estimating optimal actions through upper expectile regression on state values, suppressing extrapolation errors via asymmetric Q-updates that prioritize in-distribution high-value actions. Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) combines a variational autoencoder (VAE) with Q-filtering to restrict policy updates to data-supported regions, minimizing overestimation risks. Conservative Q-Learning (CQL) (Kumar et al., 2020) explicitly regularizes Q-values with theoretical guarantees, penalizing out-of-distribution (OOD) actions through conservative value bounds. In contrast, Decision Transformer (DT) (Chen et al., 2021) circumvents value functions entirely by reframing RL as sequence modeling, autoregressively generating actions conditioned on historical trajectories to eliminate bootstrapping errors and OOD queries.
2.2. Diffusion Model for Decision Making
@@
Diffusion models (Ho et al., 2020; Song et al., 2020), built upon stochastic denoising processes, have emerged as a dominant paradigm in image synthesis, capable of generating high-fidelity data through Markov chain-based iterative refinement. These models can be divided into two categories: conditional and unconditional paradigms. Conditional diffusion models (Fu et al., 2024) incorporate additional information (e.g., class labels, text prompts, or environmental states) to steer the generation process, enabling targeted synthesis. While unconditional diffusion models (Heng et al., 2024) generate data from pure noise without external guidance, excelling in capturing the intrinsic distribution of the training data. Owing to the powerful capabilities of diffusion models, they have been widely applied across various fields, particularly in decision-making. (Hansen-Estruch et al., 2023) proposes IDQL, an RL framework that integrates implicit Q-learning with diffusion policies in an actor-critic architecture. (Wang et al., 2022) leverages diffusion models to capture complex, multimodal behavior distributions from static datasets. (Chen et al., 2022) uses diffusion models to more accurately fit the complex distribution of behavioral policies and avoids selecting unseen actions that are out of the data distribution during training.
2.3. Real-Time Auto-Bidding
@@
Auto-Bidding in online advertising dynamically adjusts bidding strategies to maximize effectiveness (e.g., clicks, conversions, or sales) under constraints such as budgets and return on ad spend (ROAS). RL-based methods are widely used in auto-bidding because they dynamically optimize bids in complex advertising environments and balance exploration and exploitation. (Cai et al., 2017) proposes an RL-based framework to optimize real-time bidding strategies in display advertising. (He et al., 2021) proposes a generalized framework for optimizing real-time bidding strategies under diverse budget and key performance indicator (KPI) constraints, dynamically adjusting bid parameters via RL to maximize campaign value while ensuring compliance with advertiser requirements. (Jin et al., 2018) explores multi-agent RL for real-time bidding in display advertising. Recent studies (Gao et al., 2025; Guo et al., 2024b) have increasingly explored generative auto-bidding strategies in online advertising. (Gao et al., 2025) leverages Decision Transformer (Chen et al., 2021) to model sequential dependencies and historical auction dynamics, enabling adaptive bidding strategy optimization by autoregressively generating context-aware bid adjustments aligned with long-term advertiser objectives. (Guo et al., 2024b) directly models the correlation between the return and the entire trajectory and uses a conditional diffusion for bid generation.
@@ 3. Preliminary
In this section, we outline the general bidding and auction tasks framework and examine the architecture of generative diffusion models. To prevent any confusion regarding subscripts in the subsequent discussion, it should be noted that $k$ denotes the diffusion time-steps, while $t$ represents the position within the trajectory $\tau$.
3.1. Formulation of Auto-Bidding Problem
@@ 3.2. Basic Diffusion Model
3.3. Diffusion with Classifier-Free Guidance
@@
(2) $\max_{\theta}\mathbb{E}_{\tau\sim D}\left[\log p_{\theta}\left(x_{0}(\tau)\mid y(\tau)\right)\right],$
@@
where $q$ represents the forward process in which noises are gradually added to the trajectory, while $p_{\theta}$ is the reverse process, where a model is used for denoising.
We model the forward process $q\left(x_{k+1}(\tau)\mid x_{k}(\tau)\right)$ via diffusion over states:
@@
where $s_{t}$ is a one-dimensional vector, including real-time bidding context, and $x_{k}(\tau)$ is a noise sequence of states which can be represented by a two-dimensional array where the first dimension is the periods. The second dimension is the state values. Given $x_{k}(\tau)$, the forward process of the diffusion model can be represented as follows:
(5) $q\left(x_{k}(\tau)\mid x_{k-1}(\tau)\right)=\mathcal{N}\left(x_{k}(\tau);\sqrt{1-\beta_{k}}x_{k-1}(\tau),\beta_{k}\mathbf{I}\right).$
@@
when serving at period $t$, we first sample from an initial trajectory $x’_{k}(\tau)\sim\mathcal{N}(0,\mathbf{I})$ and assigns the history states $s_{0:t}$ into it. Then, we can sample predicted states with the reverse process recursively by
(8) $x’_{k-1}(\tau)=\mu_{\theta}\left(x’_{k}(\tau),y(\tau),k\right)+\sqrt{\beta_{k}}z,$
@@ 4. Methodology
In this section, we present a comprehensive description of our proposed EGDB  framework. We begin by formalizing the core mechanism of Blended-Forcing with optimal experts in Section 4.1. Next, in Section 4.2, we elaborate on the expert-guided conditional planner that translates blended signals into structured action sequences. Finally, Section 4.3 details the efficient action generation module designed for real-time execution. The overall architecture is illustrated in Figure 1.
4.1. Blended-Forcing with Optimal Expert
To address the bidding problem outlined in Equation 1, we initially aim to generate expert trajectories. To concentrate on a single bidding agent, we remove the subscript $i$, assuming all other agents are fixed. We transform the linear programming (LP) function into its dual form. Assuming the auction is truthful, the bidding formula will produce an outcome corresponding to the optimal primal solution $x_{j}^{e}$ (Aggarwal et al., 2024). For each query $j$:
(10) $\mathbf{x}^{e}(\tau)=(x_{1}^{e},\ldots,x_{T}^{e}),\quad x_{j}^{e}=\frac{1+\alpha_{c}\cdot C_{i}}{\alpha_{b}+\alpha_{c}}\cdot v_{j},$
where $\alpha_{b}$ represents the influence of budget constraints and $\alpha_{c}$ denotes the impact of constraints of cost. Then, we can refine the $\mathbf{x}^{e}(\tau)$ to fully capture the latent expert behaviors through the blended-forcing mechanism in the training phase. As shown in Figure 1, it comprises two operations: (i) the teacher-forcing technique ($TF(\cdot)$) to imitate the expert behaviors more stably, and (ii) the decode-forcing mechanism ($DF(\cdot)$) to explore more optimal expert behaviors through an extra Variational Auto-Encoder (VAE)(Kingma, 2013).
Teacher-Forcing (TF), initially introduced in Recurrent Neural Networks (RNNs) (Williams and Zipser, 1989), is a training technique where the ground-truth sequence is fed as input to the model at each time step, rather than using its previous predictions. This approach effectively mitigates error accumulation in autoregressive models by aligning training dynamics with the actual data distribution. However, a critical limitation arises during inference: the absence of ground truth forces the model to rely entirely on its own (potentially erroneous) predictions, resulting in a mismatch between the training and inference input distributions. This phenomenon, termed Exposure Bias, often results in degraded performance in real-world scenarios. Inspired by this technique, we introduce TF and use expert trajectories serving as ground truth sequences to help the model imitate expert behaviors.